00 is equal to ?
1
Expert's answer
2015-06-01T10:20:19-0400
If 0 means a fixed number, then 0^0 is assumed to be equal to 1 so that the function x^0 is continuous at x=0, because any real number raised to the power of 0 is equal to 1. If you take the limit of functions, then expression 0^0 is indeterminate form. Consider 0^0=e^(0*ln0)=e^(0*infty). Indeterminate form (0*infty) is reduced to either  0/0 or infty/infty with application of L'Hopital's rule or other method.